Quinine is a natural product extracted from the bark of the cinchona tree, which is native to South America. Quinine is used as an antimalarial agent.

When 1.68 g of quinine is dissolved in 25.0 g of cyclohexane, the freezing point of the solution is lowered by 4.31∘ C. The freezing point and Kf constant for cyclohexane can be found here.

Calculate the molar mass of quinine.Solvent Formula Kf valueNormal freezing Kb value Normal boiling point (°C) 0.00 5.49 6.59 C/m) Point (°c) 80.1 80.7 78.4 C/m H2O 1.86 5.12 cyclohexane C6H1 20.8 C2H6O1.99 0.512 2.53 2.92 1.22 water 100.00 benzene C6H6 ethanol
Answer #1

\DeltaTf = Kf*m

m is molality, Kf is freezing point depression constant,\DeltaTf is change in freezing point.

m = \Delta Tf / Kf

= 4.31 / 20.8

m = 0.2072

m = no of moles / mass of solvent in Kg

no of moles of solute = m * mass of solvent in Kg

= 0.2072 * 0.025

= 0.00518 moles

moles of solute = amount of solute / molar mass

molas mass of solute = 1.68 / 0.00518 = 324.32 g / mol

so molar mass of quinine is 324.32 g /mol